MsBrains (ms-brains):

Please help me with 5 Q's! I will medal!

MsBrains (ms-brains):

#1

OpenStudy (astrophysics):

Notice how the equation is set up "y = mx+b" form kind of right?

MsBrains (ms-brains):

Yes..?

OpenStudy (astrophysics):

So then first thing, we can graph the line, lets ignore the inequality for now.

OpenStudy (astrophysics):

I'm assuming you know how to graph equation of a line as you're at this stage right?

OpenStudy (astrophysics):



MsBrains (ms-brains):

Yeah, I think it might be A or C. Those two make the most sense

OpenStudy (astrophysics):

Well I guess all the lines are already graphed for you, but you should know how to graph a line in any case. So lets note one thing, when you have an inequality with \[\le ~~~or ~~~ \ge \] the line will be `solid` and if it's < or > it will be `dashed`

MsBrains (ms-brains):

Okay.. :)

OpenStudy (astrophysics):

And since it's \[\le \] we shade the area on the left side of the graph as you already mentioned you knew it was either A or C. S

MsBrains (ms-brains):

So it would be A then..? Because the line is solid

OpenStudy (astrophysics):

Correct

OpenStudy (astrophysics):

Please only ask one question per post as it gets convoluted. But for your next question plug in the values and if they're greater than -8 that's your solution

MsBrains (ms-brains):

It would be D then. (2, - 16) is greater then -8

OpenStudy (astrophysics):

What is the value when you plug them in

MsBrains (ms-brains):

When I plug what in? 3x+y > -8 = x > -8 -y / 3

OpenStudy (astrophysics):

The coordinates..

OpenStudy (astrophysics):

(2,-16) -> (x,y)

OpenStudy (astrophysics):

 test all of these (x,y) together